# Consistencies and inconsistencies in redshift-independent distances

### v1

@@ 4.1 NED-D
Our consistency results are summarised in Tables 1 for the standard candle distance indicators and 2 for the standard ruler and secondary indicators. We report, for each method, the number of galaxies that have at least two distance measurements using the given indicator, the number of measurement pairs for these galaxies, the total number of measurements in the NED-D catalogue from this indicator and the number of references reporting these measurements, the KLD $p$-value (with a traffic-light-style colouring indicating the severity of the inconsistency) and external consistency statistics. For self-inconsistent cases we show the $p$-value after rescaling the uncertainties to give $\Delta_{ij}$ the variance of a standard half-normal distribution in parentheses. The final column contains the number of other distance indicators with which the method in question is consistent ($p>0.05$), in four different hyphen-separated cases. The first number describes the fiducial case with no shifting of means or rescaling of uncertainties. The second number reports the result when the means of the $\mu$ distributions are made to coincide, thus eliminating discrepant central values as a cause of inconsistency. The third number corresponds to rescaling the uncertainties so that the $\Delta_{ij}$ distribution has the width of a standard half-normal by construction, eliminating discrepant uncertainties as a cause of uncertainty. The fourth number corresponds to the case with both shifted means and rescaled uncertainties, eliminating both of these as possible causes. The final number is the total number of other methods with which that method has overlap, i.e. the largest number of pairs that could possibly be consistent.
@@
Figure 2: Consistency $p$-value results for the NED-D distance indicators. The diagonal squares describe internal consistency whereas the non-diagonal squares describe external consistency with other methods. We show the cases with $p$-value higher than 0.05 (i.e. not rejecting the null hypothesis) in green, the cases with no reported $p$-value in white and the cases with a $p$-value lower than 0.05 in a logarithmic colour scale. We show cases with underestimated uncertainties with a backslash hatching, overestimated uncertainties with forward-slash hatching, and inconsistent means with circle hatching. If the cause of the inconsistency is neither one of these nor their combination we show no hatching. We call this an essential inconsistency.
@@
In Fig. 2 we summarize the results for all distance indicators. We show the $p$-values both for internal consistency along the diagonal and external consistency on the off-diagonal. Cases where no consistency test is possible due to no overlap in measurements are shown as empty squares. Green indicates consistency ($p>0.05$), while $p$-values below 0.05 are shown by the colour-bar, with darker red indicating lower $p$. As we use 10,000 mocks the $p$-value is calculated in units of 0.0001; anything below this (i.e. no mock KLD as large as that of the data) is shown in dark red. The hatching indicates the cause of inconsistency. Over- and underestimated uncertainties are shown by forward and backward slanting diagonal lines respectively, while inconsistent means are shown by the circle pattern. Note that some comparisons exhibit both problems and hence have both types of hatching. If neither shifting the means nor rescaling the uncertainties removes the inconsistency we label it “essentially”, indicated by no hatching.
Figure 2 contains all the information for the consistency of the distance indicators. The total number of consistent cases is 427. 219 are inconsistent due to have uncertainties that are either too large or too small (mostly too small). Six have inconsistent means, and 169 have both problematic means and uncertainties. For problems that our tests cannot detect such as combinations of the previous ones along with e.g. outliers or non-Gaussian behaviour, we show no hatching. There are 169 such cases. Finally, 1994 pairs have no shared galaxies and hence no $p$-value. Therefore, the analyses showing consistency are 427 out of 932, representing 41.8% of all cases. Combining consistent cases with those where we have identified the problem yields 821 cases, representing 88.1% of the total. Hence, even though most distance indicators are internally inconsistent, close to half of the cross-analyses are consistent, while around just one tenth are inconsistent with a non-obvious problem.
@@
N3254 2628 0.0778
N1559 15576 0.042 (0.587) $\uparrow$
N3447 1431 0.0242 (0.2722) $\uparrow$
N5584 31375 0.0196 (0.851) $\uparrow$
N0976 28 0.0166 (0.4116) $\downarrow$
N7250 465 0.0154 (0.6016) $\downarrow$
@@ 4.2 SH0ES
Fig. 5 and the top row of Table 3 indicate that the SH0ES 2022 catalogue agrees well with the null hypothesis of statistical consistency. The two galaxies used for the first rung that calibrates the PL relation are also consistent with the null hypothesis. It is particularly important for these to be consistent, as inconsistency there would propagate into all the other distances and hence the constraint on e.g. $H_{0}$. From the 38 galaxies in the second rung (M31 does not host a SNIa but it is included in the SH0ES catalogue), 31 are consistent while the remaining 7 are inconsistent due to underestimated uncertainties. The inconsistency is in most cases weak; had we take a tighter 0.01 threshold for the $p$-value, only one galaxy would have been inconsistent (N4038). Most of the inconsistent galaxies have $<$2000 pairs (apart from N1559 and N5584) and hence play a relatively small role in the consistency of the whole catalogue and the calibration of $M^{0}_{B}$ and hence $H_{0}$. Four of the seven inconsistent galaxies have underestimated uncertainties (N0976, N7250, M31 and N4038) whereas N1559, N3447 and N5584 have overestimated uncertainties. Interestingly, the inconsistent galaxies having more than 15000 measurement pairs have overestimated uncertainties, which has the effect of overestimating the uncertainty on $H_{0}$. We highlight two galaxies in lighter orange (N3982 and N0691) that would be inconsistent under a two-tailed test due to having $p>0.975$. This is due to the KLD being surprisingly low relative to the mock data distribution, indicating that the data conforms more closely to a standard half-normal than would be expected under the assumption of consistency given the uncertainties.
@@ 5.2 Comparison with the literature
Our NED-D analysis builds upon Singh (2021), where possible evidence for systematic underestimation of distance uncertainties was found. This could imply strong systematics on the late-time measurements of $H_{0}$, for example (Freedman et al., 2019a; Riess et al., 2022; Anand et al., 2022; Scolnic et al., 2023; Riess & Breuval, 2024), which might overestimate the tension with Planck. However, as mentioned before, Singh (2021) uses a non-symmetric expression for the discrepancy between two measurements that does not include the uncertainty of both measurements. This is in contrast to the Mahalanobis distance (Eq. 3) which takes both of them into account. The non-symmetric expression gives higher sigma differences as one of the uncertainties is neglected, artificially increasing discrepancies. Even so, in agreement with  Singh (2021) we infer inconsistencies in most of the NED-D methods.
@@
Nevertheless there is a plethora of possible problems with the Cepheid distance ladder (Owens et al., 2021, 2022; Madore & Freedman, 2024b; Owens et al., 2024; Madore & Freedman, 2024a), as well as results from the Chicago-Carnegie Hubble Program (CCHP) using new data from the James Webb Space Telescope (JWST) indicating that $H_{0}\approx 70$ km/s/Mpc with an insignificant tension with Planck (Freedman et al., 2024). This agrees with earlier CCHP results based on TRGB (Freedman et al., 2019a), and includes distances measured by the novel J-region Asymptotic Giant Branch (JAGB) method which boasts a number of advantages over Cepheids and TRGB (Madore & Freedman, 2020; Lee et al., 2024).
@@ 6 Conclusions
We study the consistency of 66 different methods for estimating redshift-independent distances from the NASA/IPAC Extragalactic Database of Distances (NED-D; Steer et al. 2016) and SH0ES 2022 (Riess et al., 2022) catalogues. This is achieved by selecting galaxies that have at least two distance measurements by a given method (“internal consistency”) or at least one measurement by each of two different methods (“external consistency”). We use the Mahalanobis distance to compute the distance differences in units of sigma, forming a distribution for each consistency test. If the distance estimates are consistent (scattered from a common value by their quoted uncertainties), this distribution should form a standard half-normal. To test if this is the case, we compute the Kullback–Leibler (KL) divergence of the empirical distribution and of 10,000 realizations of random samples under the null hypothesis of consistency and according to the measurements’ uncertainties. We compute the $p$-value as the fraction of realisations with a KL divergence higher than the value in the real data, adopting a threshold of 0.05 to discard the null-hypothesis of consistency. This indicates under- or overestimated uncertainties, inconsistent means, statistical outliers, non-Gaussian uncertainties and/or systematic errors. We perform a similar analysis for the SH0ES 2022 catalogue.
@@
We also calculate the average measured distance modulus of each indicator across all galaxies relative to the mean of all indicators. We find that from the best set of indicators, Proper Motion is the only one with a considerable discrepancy in average returned distance (17.1% underestimation) relative to the mean. Outside the set of best indicators, some overestimate the luminosity distances by more than 95%; these are SNIa SDSS, H I + optical distribution and Wolf-Rayet. AGN time lag shows an overestimation higher than 25% and Mass model an underestimation higher than 20%. This indicates that some distance methods in NED-D are strongly affected by systematics and thus unsuitable for cosmological studies. We perform 932 analyses in total, including both internal and external consistency, of which 427 are consistent (41.8% of the total and 12 of them are from internally consistency analyses and the remainder externally consistency analyses). We repeat the analyses for the inconsistency cases by shifting the means and rescaling the uncertainties and find that 394 cases are consistent after these modifications. In the remaining $\sim$1/10 of cases the inconsistency is not due just to incorrect central values or uncertainties and we cannot identify the cause.
@@
In summary, the NED-D distance indicator measurements are subject to problems like systematics, outliers, non-Gaussian behaviour and misestimated uncertainties. This is perhaps unsurprising given the heterogeneity of the measurements included, going back $\sim$45 years. Thus the full catalogue is not suitable for construction of a robust distance ladder. Even so, several papers have used the NED-D data for studies in cosmology (Freedman et al., 2019b; Chaparro-Molano et al., 2019; Steer, 2020; Zaninetti, 2023). This shows that by selecting a reliable subset of distances, it is possible to determine galaxies’ distances accurately. Some of the inconsistency comes from non-independent and unreliable distance measurements that were impossible to exclude in the search procedure used to construct the catalogue. By removing these problematic data, the overall consistency might be considerably enhanced, enabling cosmological analysis; we will report elsewhere on our attempts in this direction. The more homogeneous and quality-controlled SH0ES 2022 catalogue, however, has no significant evident problems. We hope the framework devised here will contribute to the future construction of a multiply cross-checked and fully robust distance ladder employing a range of methods.
@@
The NED-D data is available at https://ned.ipac.caltech.edu/Library/Distances/, and the SH0ES 2022 data at https://github.com/PantheonPlusSH0ES/DataRelease. We will make our code publicly available on publication of the paper. Other data underlying the article may be made available upon request to the authors.

### v2

@@ 4.1 NED-D
Our consistency results are summarised in Tables 1 for the standard candle distance indicators and 2 for the standard ruler and secondary indicators. We report, for each method, the number of galaxies that have at least two distance measurements using the given indicator, the number of measurement pairs for these galaxies, the total number of measurements in the NED-D catalogue from this indicator and the number of references reporting these measurements, the KLD $p$-value (with a traffic-light-style colouring indicating the severity of the inconsistency) and external consistency statistics. For self-inconsistent cases we show the $p$-value after rescaling the uncertainties to give $\Delta_{ij}$ the variance of a standard half-normal distribution in parentheses. The final column contains the number of other distance indicators with which the method in question is consistent ($p>0.05$), in five different hyphen-separated cases. The first number describes the fiducial case with no shifting of means or rescaling of uncertainties. The second number gives the number of pairs that are made consistent when the means of the $\mu$ distributions are made to coincide, thus eliminating discrepant central values as a cause of inconsistency. The third number is the same but for rescaling the uncertainties so that the $\Delta_{ij}$ distribution has the width of a standard half-normal by construction, eliminating discrepant uncertainties as a cause of uncertainty. The fourth number corresponds to the case in which either shifting the means or rescaled the uncertainties produces consistency, and the fifth to the case where doing both is required to produce consistency. The final number is the total number of other methods with which that method has overlap, i.e. the largest number of pairs that could possibly be consistent.
@@
Figure 2: Consistency $p$-value results for the NED-D distance indicators. The diagonal squares describe internal consistency whereas the non-diagonal squares describe external consistency with other methods. We show the cases with $p$-value higher than 0.05 (i.e. not rejecting the null hypothesis) in green, the cases with no reported $p$-value in white and the cases with a $p$-value lower than 0.05 in a logarithmic colour scale. We show cases with underestimated uncertainties with a backslash hatching, overestimated uncertainties with forward-slash hatching, and inconsistent means with circle hatching. If either shifting the means or rescaling the uncertainties produces consistency we show both types of hatching, while if both are required we show star hatching. If the cause of the inconsistency is neither one of these nor their combination (i.e. it remains inconsistent after shifting the means and rescaling the uncertainties) we show no hatching. We call this an essential inconsistency.
@@
In Fig. 2 we summarize the results for all distance indicators. We show the $p$-values both for internal consistency along the diagonal and external consistency on the off-diagonal. Cases where no consistency test is possible due to no overlap in measurements are shown as empty squares. Green indicates consistency ($p>0.05$), while $p$-values below 0.05 are shown by the colour-bar, with darker red indicating lower $p$. As we use 10,000 mocks the $p$-value is calculated in units of 0.0001; anything below this (i.e. no mock KLD as large as that of the data) is shown in dark red. The hatching indicates the cause of inconsistency. Over- and underestimated uncertainties are shown by forward and backward slanting diagonal lines respectively, while inconsistent means are shown by the circle pattern. Both types of hatching indicate that both problems are present to some degree, such that resolving either one produces consistency, and star hatching that both problems are present to a greater degree such that resolving both is required to remove the inconsistency. If neither shifting the means nor rescaling the uncertainties removes the inconsistency we label it “essential”, indicated by no hatching.
Figure 2 contains all the information for the consistency of the distance indicators. The total number of consistent pairs is 427 (41.8% of all pairs). 221 are inconsistent due to have uncertainties that are either too large or too small (mostly too small). Seven have inconsistent means, 109 are made consistent by matching either the means or the uncertainties, and five only by matching both. For problems that our tests cannot detect such as combinations of the previous ones along with e.g. outliers or non-Gaussian behaviour, we show no hatching. There are 163 such cases. Finally, 1994 pairs have no shared galaxies and hence no $p$-value. (Note that each pair is counted twice in Tables 1 and 2, so that the sums of the corresponding columns are twice the numbers quoted above.) Combining a priori consistent cases with those where we have identified the problem yields 769 cases, representing 82.5% of the total. Hence, even though most distance indicators are internally inconsistent, close to half of the cross-analyses are consistent, while only around a fifth are inconsistent due to an undiagnosed problem.
@@
N3370 2628 0.0778
N2442 15576 0.042 (0.587) $\uparrow$
N3583 1431 0.0242 (0.2722) $\uparrow$
N5643 31375 0.0196 (0.851) $\uparrow$
N0105 28 0.0166 (0.4116) $\downarrow$
N7329 465 0.0154 (0.6016) $\downarrow$
@@ 4.2 SH0ES
Fig. 5 and the top row of Table 3 indicate that the SH0ES 2022 catalogue agrees well with the null hypothesis of statistical consistency. The two galaxies used for the first rung that calibrates the PL relation are also consistent with the null hypothesis. It is particularly important for these to be consistent, as inconsistency there would propagate into all the other distances and hence the constraint on e.g. $H_{0}$. From the 38 galaxies in the second rung (M31 does not host a SNIa but it is included in the SH0ES catalogue), 31 are consistent while the remaining 7 are inconsistent due to underestimated uncertainties. Note that most of the $p$-values only narrowly pass the threshold for rejection of the null hypothesis of consistency; had we take a tighter 0.01 threshold for the $p$-value, only one galaxy would have been considered inconsistent (N4424). This contrasts with NED-D where the great majority of inconsistent cases have $p<0.01$. Most of the inconsistent galaxies have $<$2000 pairs (apart from N2442 and N5643) and hence play a relatively small role in the consistency of the whole catalogue and the calibration of $M^{0}_{B}$ and hence $H_{0}$. Four of the seven inconsistent galaxies have underestimated uncertainties (N0105, N7329, M31 and N4424) whereas N2442, N3583 and N5643 have overestimated uncertainties. Interestingly, the inconsistent galaxies having more than 15000 measurement pairs have overestimated uncertainties, which has the effect of overestimating the uncertainty on $H_{0}$. We highlight two galaxies in lighter orange (N4038 and N1015) that would be inconsistent under a two-tailed test due to having $p>0.975$. This is due to the KLD being surprisingly low relative to the mock data distribution, indicating that the data conforms more closely to a standard half-normal than would be expected under the assumption of consistency given the uncertainties.
@@ 5.2 Comparison with the literature
Our NED-D analysis builds upon Singh (2021), where possible evidence for systematic underestimation of distance uncertainties was found. This could imply strong systematics on the late-time measurements of $H_{0}$, for example (Freedman et al., 2019a; Riess et al., 2022; Anand et al., 2022; Scolnic et al., 2023; Riess & Breuval, 2024), which might overestimate the tension with Planck. Despite using a different metric to quantify tension, and a considerably more sophisticated statistical framework, we agree with  Singh (2021) that most NED-D methods are inconsistent.
@@
Nevertheless there is a plethora of possible problems with the Cepheid distance ladder (Owens et al., 2021, 2022; Madore & Freedman, 2024b; Owens et al., 2024; Madore & Freedman, 2024a), as well as results from the Chicago-Carnegie Hubble Program (CCHP) using new data from the James Webb Space Telescope (JWST) indicating that $H_{0}\approx 70$ km/s/Mpc with an insignificant tension with Planck (Freedman et al., 2024; Hoyt et al., 2025). This agrees with earlier CCHP results based on TRGB (Freedman et al., 2019a), and includes distances measured by the novel J-region Asymptotic Giant Branch (JAGB) method which boasts a number of advantages over Cepheids and TRGB (Madore & Freedman, 2020; Lee et al., 2024).
@@ 6 Conclusions
We study the consistency of 66 different methods for estimating redshift-independent distances from the NASA/IPAC Extragalactic Database of Distances (NED-D; Steer et al. 2016) and SH0ES 2022 (Riess et al., 2022) catalogues. This is achieved by selecting galaxies that have at least two distance measurements by a given method (“internal consistency”) or at least one measurement by each of two different methods (“external consistency”). We use the Mahalanobis distance to compute the distance differences in units of sigma, forming a distribution for each consistency test. If the distance estimates are consistent (scattered from a common value by their quoted uncertainties), this distribution should form a standard half-normal. To test if this is the case, we compute the Kullback–Leibler (KL) divergence of the empirical distribution and of 10,000 realizations of random samples under the null hypothesis of consistency and according to the measurements’ uncertainties. We compute the $p$-value as the fraction of realisations with a KL divergence higher than the value in the real data, adopting a threshold of 0.05 to discard the null hypothesis of consistency. This indicates significant evidence for under- or overestimated uncertainties, inconsistent means, statistical outliers, non-Gaussian uncertainties and/or systematic errors. When $p>0.05$, there is not enough evidence to discard the null hypothesis, which we denote by “consistent.” We perform a similar analysis for the SH0ES 2022 catalogue.
@@
We also calculate the average measured distance modulus of each indicator across all galaxies relative to the mean of all indicators. We find that from the best set of indicators, Proper Motion is the only one with a considerable discrepancy in average returned distance (17.1% underestimation) relative to the mean. Outside the set of best indicators, some overestimate the luminosity distances by more than 95%; these are SNIa SDSS, H I + optical distribution and Wolf-Rayet. AGN time lag shows an overestimation higher than 25% and Mass model an underestimation higher than 20%. This indicates that some distance methods in NED-D are strongly affected by systematics and thus unsuitable for cosmological studies. We perform 932 analyses in total, including both internal and external consistency, of which 427 are consistent (41.8% of the total and 12 of them are from internally consistency analyses and the remainder externally consistency analyses). We repeat the analyses for the inconsistency cases by shifting the means and rescaling the uncertainties and find that 342 cases are made consistent by one or more of these modifications. In the remaining $\sim$1/5 of cases the inconsistency is not due only to incorrect central values or uncertainties, and we cannot identify the cause.
@@
In summary, the NED-D distance indicator measurements are subject to problems such as systematics, outliers, non-Gaussian behaviour and misestimated uncertainties. This is perhaps unsurprising given the heterogeneity of the measurements included, going back $\sim$45 years. Thus the full catalogue is not suitable for construction of a robust distance ladder. Even so, several papers have used the NED-D data for studies in cosmology (Freedman et al., 2019b; Chaparro-Molano et al., 2019; Steer, 2020; Zaninetti, 2023). By selecting a reliable subset of distances, it is however possible to determine galaxies’ distances accurately. Some of the inconsistency comes from non-independent and unreliable distance measurements that were impossible to exclude in the search procedure used to construct the catalogue. By removing these problematic data, the overall consistency might be considerably enhanced, enabling cosmological analysis; we will report elsewhere on our attempts in this direction. The more homogeneous and quality-controlled SH0ES 2022 catalogue, however, has no significant evident problems. The framework we develop may be simply applied to any catalogue of distances (or other quantities) to assess their consistency; we therefore hope that it will contribute to the future construction of a multiply cross-checked and fully robust distance ladder employing a wide range of methods.
@@
The code required to reproduce the results in this article are available at https://github.com/antonionajeraq/consistency_redshift_independent_galaxies.git. The NED-D data is available at https://ned.ipac.caltech.edu/Library/Distances/, and the SH0ES 2022 data at https://github.com/PantheonPlusSH0ES/DataRelease.